A baseball is dropped from a stadium seat that is 81 feet above the ground. Its height s in feet after t seconds is given by ​s(t)=81-16t^2. Estimate to the nearest tenth of a second how long it takes for the baseball to strike the ground.
​(Simplify your answer. Round to the nearest tenth as​ needed.)
thanks!(:

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A baseball is dropped from a stadium seat that is 81 feet above the ground. Its height s in feet after t seconds is given by ​s(t)=81-16t^2. Estimate to the nearest tenth of a second how long it takes for the baseball to strike the ground.
=======================================
​s(t)=81-16t^2 = 0
Solve for t, ignore the negative value.

Answer by ikleyn(52781)

.

As different English words are combined in this post, they don't add up to meaningful text.